# Large Language Model-based Data Science Agent: A Survey

### v1

@@ 5 Benchmark
Table 4 presents a summary of curated benchmarks designed for evaluating LLM-based agents from a data science perspective. Notable examples include ML-Bench Tang et al. (2023) for multimodal and time-series tasks and DSBench Jing et al. (2024) for data analysis and modeling. Specialized benchmarks such as GeoAgent-Bench Liao et al. (2024) provide evaluation frameworks for geospatial analysis, while MatPlotAgent-Bench Yang et al. (2024b) focuses on visualization.
Beyond general data science, several benchmarks target specialized domains. FoodPuzzle Huang et al. (2024b) explores computational approaches to understanding flavor profiles at the molecular level. Merrill et al. (2024) evaluates LLM capabilities in personal health reasoning with wearable data. AgentClinic Schmidgall et al. (2024) assesses patient interactions and multimodal clinical data processing. GenoTEX Liu & Wang (2024) benchmarks LLMs in genomics, covering dataset selection, preprocessing, and statistical analysis.
Additionally, TheAgentCompany Xu et al. (2024) serves as a practical evaluation framework for AI-driven automation in corporate environments, covering tasks in data science, software engineering, and business operations. These benchmarks provide a foundational dataset and evaluation standard for LLM-based agent systems.
@@
Benchmark Source # of Tasks Task Types BLADE Gu et al. (2024) Literature 714 multiple-choice and ground truth-analysis studies InfiAgent-DABench Hu et al. (2024a) GitHub 257 Data Analysis and ML ML-Bench Tang et al. (2023) GitHub 9641(168) Multimodal, time series, audio, LLM, vision, biomedical, graphs DevBench Li et al. (2024a) Github 22 DL, CV, and NLP SUPER Bogin et al. (2024) GitHub 799 Research data science challenges DA-Code Huang et al. (2024c) Kaggle, GitHub 500 Data wrangling, EDA, ML FeatEng Pietruszka et al. (2024) Kaggle 103 Classification, regression, feature engineering Tapilot Li et al. (2024e) Kaggle 1024 Data analysis, request clarification MLE-Bench Chan et al. (2024) Kaggle 75 Modelling Tasks(image, video, LLMs, tabular) DSBench Jing et al. (2024) ModelOff, Kaggle 540 Data analysis, modeling PyBench Zhang et al. (2024c) Kaggle, Arxiv, multimedia files 143 Data Analysis, ML, image, text, and audio analysis DSEval Zhang et al. (2024d) Tutorials, StackOverflow, Kaggle 825 Data analysis Spider2-V Cao et al. (2024) Tutorials, enterprise applications 494 Warehousing, transformation, visualization MatPlotAgent-Bench Yang et al. (2024b) Matplotlib, OriginLab 100 Standard and advanced visualization ScienceAgentBench Chen et al. (2024c) Peer-reviewed publications 102 Data processing, modeling, visualization SagePilot Liao et al. (2024) External datasets 276 SQL-related tasks Text2Analysis He et al. (2024) LLM and human-generated 2249 Data analysis, modeling DataNarrative Islam et al. (2024b) Pew Research, Tableau Public, GapMinder 1449 Data-driven visualization, storytelling Insigt-Bench Sahu et al. (2024) ServiceNow platform 100 Data Analysis tasks MMAU Yin et al. (2024) In-house, Kaggle, DeepMind-Math 20 DS/ML, contest-level coding, math GeoAgent-Bench Liao et al. (2024) GitHub, tutorials, LLM generation 19,504 Single-turn, multi-turn in Geo-spatial analysis MLAgentBench Huang et al. (2023) Kaggle, canonical datasets 13 Research Image, text, graph, tabular, time series ML tasks Merrill et al. (2024) Human experts, wearable data 4172 Numerical, open-ended reasoning in personal health AgentClinic Schmidgall et al. (2024) USMLE, MIMICIV, NEJM, MedAQ 535 Patient interaction, multimodal data collection in clinics GenoTEX Liu & Wang (2024) GEO, TCGA, NCBI Gene Database 1146 Dataset selection, processioning, statistical analysis in Genomics TheAgentCompany Xu et al. (2024) Company websites, human examples 175 DS tasks in company settings FoodPuzzle Huang et al. (2024b) FlavorDB 2744 Molecular prediction and profile completion in flavor science MLGym Nathani et al. (2025) Publications and datasets 13 DS, CV, NLP, RL, and game theory DataSciBench Zhang et al. (2025a) CodeGeeX, BCB, and human 519 Data Analysis, modeling, data Visualization Chen et al. (2024a) CREEDS and GEO GEO Database and Drug Repurposing Database in medical fields BIODSA-1K Wang et al. (2025b) Publications 1029 Biomedical hypothesis and analysis DS-1000Lai et al. (2023) Stackoverflow 1000 Code generation in DS BioDSBench Wang et al. (2024b) Published studies, TCGA-type genomics, and clinical data 293 Biomedical coding tasks ML-Dev-Bench Padigela et al. (2025) Unknown 30 Data processing, modeling, API integration TimeSeriesGym Cai et al. (2025) Kaggle, Github, publications, hand-crafted 34 Time-series problem
@@ 6 Future Research Opportunity
In this section, we outline future directions inspired by recent advancements and existing challenges identified to stimulate subsequent research in LLM-based data science agents.
6.1 Trainable architecture
Inspired by pioneering works such as EvoMACHu et al. (2024b), there is substantial potential in exploring dynamically refine agent architecture. This approach facilitates continuous optimization of system efficacy across varying data science domains. Emerging methodologiesYuksekgonul et al. (2024); Hu et al. (2025) propose leveraging backpropagation-inspired textual gradients to systematically adjust and improve agent architectures. Future research might investigate automated structural generation methods capable of autonomously deriving optimal agent architectures tailored explicitly to domain-specific requirements, significantly enhancing both scalability and adaptability of LLM-based agents.
6.2 Advanced Reflection Mechanisms
Current reflection strategies predominantly address short-term corrections through immediate feedback mechanisms. Expanding these mechanisms to include comprehensive long-term reflective processes, such as history-window-based learning and unsupervised self-correction, presents significant opportunities. Implementing iterative refinement methodologies and advanced error-tracing capabilities can substantially enhance the robustness and reliability of agents over extended operational periods. Furthermore, an in-depth examination of the underlying reasons behind reflective failuresCemri et al. (2025) could provide crucial insights. Such research may lead to the development of sophisticated reflective frameworks capable of proactively identifying and mitigating systematic inefficiencies.
6.3 Multimodal Processing
Integrating vision-language models (VLMs) with LLM-based agents offers a substantial opportunity to enhance their interpretative capabilities, particularly in analyzing visual data prevalent in statistical and analytical reports. The incorporation of specialized multimodal agents demonstrates a clear potential for effectively interpreting visual information, generating relevant feedback, and correcting inaccuraciesYang et al. (2024b). Future research may explore hybrid models that fuse textual and visual modalities more seamlessly, leveraging cross-modal embeddings to improve the precision of visual data interpretation. Additionally, developing specialized visual-feedback agents capable of autonomously interpreting and refining graphical data could significantly improve the accuracy, reliability, and utility of insights derived from visual analytics within complex data-driven environments.

### v2

@@ 5 Benchmark
A wide range of benchmarks has been developed to evaluate LLM-based data science agents, each reflecting different assumptions about what “data science competence” entails. Table 5 summarizes the major suites. General-purpose benchmarks such as ML-Bench tang2023ml and DSBench jing2024dsbench assess core analytical skills across preprocessing, modeling, feature engineering, and multimodal reasoning. Other benchmarks specialize in particular workflow stages, such as visualization in MatPlotAgent-Bench yang2024matplotagent or geospatial analysis in GeoAgent-Bench liao2024towards. Domain-focused benchmarks, such as FoodPuzzle huang2024foodpuzzle, GenoTEX liu2024genotex, or AgentClinic schmidgall2024agentclinic, measure scientific and data-driven reasoning under discipline-specific constraints. In addition, TheAgentCompany xu2024theagentcompany provides a practical evaluation framework covering end-to-end automation tasks in corporate environments, including data science, engineering, and business operations.
While these benchmarks provide broad coverage and standardized evaluation settings, they also exhibit structural limitations that may mischaracterize an agent’s true capabilities. First, most datasets evaluate short, isolated tasks. Suites such as DSBench or ML-Bench typically test a single operation (e.g., computing a statistic, performing a small transformation, or fitting a lightweight model), thus missing the errors accumulated across multi-stage workflows. Second, rigid task formats allow models to exploit surface regularities rather than reasoning. In MatPlotAgent-Bench, for example, many tasks share highly similar visualization templates (e.g., generating a 2×3 grid of boxplots), so an agent may match the template instead of demonstrating genuine analytical understanding and reasoning. Third, benchmark inputs are often clean, complete, and unambiguous. But LLM agents frequently fail when inputs are messy or underspecified (e.g., irregular schemas or missing values), a scenario that current benchmarks like GeoAgent-Bench still rarely capture.
Each benchmark therefore captures only a particular slice of the capabilities required for real data science practice. General-purpose suites test broad coverage but not long-horizon robustness; workflow-specific and domain-specific benchmarks evaluate deeper skills but in narrower settings. A comprehensive evaluation therefore requires viewing performance across multiple benchmarks, each revealing different aspects of reliability, reasoning, and robustness.
@@
Benchmark Source # of Tasks Task Types BLADE gu2024blade Literature 714 multiple-choice and ground truth-analysis studies InfiAgent-DABench hu2024infiagent GitHub 257 Data Analysis and ML ML-Bench tang2023ml GitHub 9641(168) Multimodal, time series, audio, LLM, vision, biomedical, graphs DevBench li2024devbench Github 22 DL, CV, and NLP SUPER bogin2024super GitHub 799 Research data science challenges DA-Code huang2024code Kaggle, GitHub 500 Data wrangling, EDA, ML FeatEng pietruszka2024can Kaggle 103 Classification, regression, feature engineering Tapilot li2024tapilot Kaggle 1024 Data analysis, request clarification MLE-Bench chan2024mle Kaggle 75 Modelling Tasks(image, video, LLMs, tabular) DSBench jing2024dsbench ModelOff, Kaggle 540 Data analysis, modeling PyBench zhang2024pybench Kaggle, Arxiv, multimedia files 143 Data Analysis, ML, image, text, and audio analysis DSEval zhang2024benchmarking Tutorials, StackOverflow, Kaggle 825 Data analysis Spider2-V cao2024spider2 Tutorials, enterprise applications 494 Warehousing, transformation, visualization MatPlotAgent-Bench yang2024matplotagent Matplotlib, OriginLab 100 Standard and advanced visualization ScienceAgentBench chen2024scienceagentbench Peer-reviewed publications 102 Data processing, modeling, visualization SagePilot liao2024towards External datasets 276 SQL-related tasks Text2Analysis he2024text2analysis LLM and human-generated 2249 Data analysis, modeling DataNarrative islam2024datanarrative Pew Research, Tableau Public, GapMinder 1449 Data-driven visualization, storytelling Insigt-Bench sahu2024insightbench ServiceNow platform 100 Data Analysis tasks MMAU yin2024mmau In-house, Kaggle, DeepMind-Math 20 DS/ML, contest-level coding, math GeoAgent-Bench liao2024towards GitHub, tutorials, LLM generation 19,504 Single-turn, multi-turn in Geo-spatial analysis MLAgentBench huang2023benchmarking Kaggle, canonical datasets 13 Research Image, text, graph, tabular, time series ML tasks merrill2024transforming Human experts, wearable data 4172 Numerical, open-ended reasoning in personal health AgentClinic schmidgall2024agentclinic USMLE, MIMICIV, NEJM, MedAQ 535 Patient interaction, multimodal data collection in clinics GenoTEX liu2024genotex GEO, TCGA, NCBI Gene Database 1146 Dataset selection, processioning, statistical analysis in Genomics TheAgentCompany xu2024theagentcompany Company websites, human examples 175 DS tasks in company settings FoodPuzzle huang2024foodpuzzle FlavorDB 2744 Molecular prediction and profile completion in flavor science MLGym nathani2025mlgym Publications and datasets 13 DS, CV, NLP, RL, and game theory DataSciBench zhang2025datascibench CodeGeeX, BCB, and human 519 Data Analysis, modeling, data Visualization chen2024multi CREEDS and GEO GEO Database and Drug Repurposing Database in medical fields BIODSA-1K wang2025biodsa Publications 1029 Biomedical hypothesis and analysis DS-1000lai2023ds Stackoverflow 1000 Code generation in DS BioDSBench wang2024can Published studies, TCGA-type genomics, and clinical data 293 Biomedical coding tasks ML-Dev-Bench padigela2025ml Unknown 30 Data processing, modeling, API integration TimeSeriesGym cai2025timeseriesgym Kaggle, Github, publications, hand-crafted 34 Time-series problem
@@ 6 Future Research Opportunity
Building on the limitations discussed, several research opportunities emerge for improving the reliability of LLM-based data science agents. The core challenges of current systems lie in their vulnerability to hallucination, tendency to generate fragile and non-reproducible code, difficulty maintaining context over long-horizon workflows, and reliance on rigid agent pipelines. The following subsections outline three promising directions, each targeting a different aspect of these limitations and suggesting how next-generation agent systems may overcome them.
6.1 Data-Centric Diagnostics
As discussed earlier, many high-impact failures in LLM-based data science agents originate not from the agent’s reasoning, but from properties of the data itself. Agents frequently misinterpret irregular schemas, treat categorical columns as numerical, overlook missing-value structures that distort statistical estimates, or apply methods incompatible with the underlying distribution. Current reflection mechanisms are blind to such data issues: they only inspect the agent’s generated text or code, leaving foundational data errors undetected. This limitation explains why agents often produce “correctly reasoned” but fundamentally invalid analytical results. Future research should therefore develop data-centric diagnostic modules that automatically evaluate schema consistency, detect anomalous or distribution-shifted subsets, identify type or semantic mismatches, and trace how specific data defects influence downstream modeling outcomes. By elevating data quality analysis to a first-class reflective capability, agents would be better equipped to prevent systematic failure modes that current systems cannot even observe.
6.2 Uncertainty-Aware Workflow Planning
Another highlighted limitation is that current LLM-based agents operate without any awareness of uncertainty. When the model is unsure about a column’s meaning, a user’s intent, or the correctness of an intermediate step, its downstream behavior does not change and the uncertain output carries the same operational consequences as a confident one. Because agents lack mechanisms to detect low-confidence reasoning, they do not pause execution to request clarification, reconsider assumptions, expand the search space, or branch the workflow. As a result, ambiguous or underspecified steps often trigger confident hallucinations that steer the entire pipeline in an incorrect direction, with errors compounding across subsequent stages. Future research may therefore focus on developing uncertainty-aware workflow planning, where agents explicitly model their confidence and adapt their behavior accordingly. Such systems could flag uncertain decisions, trigger query refinement, invoke stronger verification modes, or explore multiple candidate workflows when ambiguity is high. Introducing uncertainty as a first-class signal would allow agents to dynamically modulate autonomy and mitigate the error-propagation patterns that are pervasive in long-horizon data science tasks.
6.3 Pipeline-Level Reflection
A further limitation is that existing reflection mechanisms operate only at the step level. Reviewers check the latest code snippet, self-reflection revises the previous response, and unit tests validate single functions. Yet data science workflows are inherently multi-stage, and small upstream issues, such as incorrect filtering, misaligned merges, omitted preprocessing, often propagate silently into feature engineering, model training, and evaluation. Current agents lack the ability to revisit earlier stages, attribute downstream errors to upstream decisions, or assess the global coherence of the entire analytical pipeline. Future research may address this gap by developing pipeline-level reflection: constructing explicit workflow graphs, tracing how information and errors flow across steps, identifying high-risk nodes, performing stage-level sanity checks, and enabling global revisions that correct the pipeline holistically rather than locally. Such mechanisms would transform reflection from shallow post-hoc editing into a principled, system-level reliability process.